# Choked precessing jets in tidal disruption events and high-energy neutrinos

### v1

@@ 2 Jet propagation and the break-out condition for precessing jets
After the stellar disruption, an accretion disk is formed in the orbital plane of the star whose normal is generally misaligned by an angle $\theta_{\rm LS}$ with respect to the SMBH spin axis. The accretion disk undergoes Lense-Thirring precession around the black hole spin axis and the jet is aligned with the instantaneous disk angular momentum (Teboul & Metzger, 2023; Lu et al., 2023). Matter falls back to the disk at a highly super-Eddington accretion rate $\dot{M}$, driving a disk wind of velocity $\beta_{\rm w}=v_{\rm w}/c$ and mass-loss rate $\dot{M}_{\rm w}=f_{\rm w}\dot{M}$, where $f_{\rm w}<1$ is the fraction of the mass fallback rate lost in the wind. Metzger & Stone (2016) argued that only a small fraction the returning debris ends up reaching the inner circular disk and accreting onto the SMBH, with the vast majority becoming unbound in a wind outflow. Due to the rapid precession, these disk winds will inevitably interact and collide with each other on large scales, leading to a quasi-spherical outflow.
We consider a relativistic jet with a kinetic luminosity $L_{\rm j}$, an initial half-opening angle $\theta_{0}$ and a bulk Lorentz factor $\Gamma=1/\theta_{0}$. As the jet advances in the quasi-spherical wind envelope, the jet drives a bow shock ahead of it. The jet is capped by a termination shock, and a reverse shock propagates back into the jet, where the jet is decelerated and heated. The hot head material spills sideways, forming a cocoon that engulfs the jet and collimate it. The velocity of the jet head is determined by the balance of pressure between ambient gas and the jet head (Matzner, 2003). In a moving wind envelope, we can express the velocity in the frame of the wind envelope by a coordinate transformation (e.g., Ioka & Nakamura, 2018; Hamidani et al., 2020; Gottlieb & Nakar, 2022; Lu et al., 2023):
@@
$\tilde{L}=\frac{L_{\rm j}}{\Sigma_{\rm j}\rho_{\rm w}(r_{\rm h})c^{3}}=\frac{L_{\rm j,iso}\beta_{\rm w}}{\dot{M}_{\rm w}c^{2}}.$ (2)
Here $\beta_{\rm h}$ is the jet head velocity in unit of the speed of light (measured in the lab frame), $r_{\rm h}$ is the radius of the jet head, $\rho_{\rm w}$ is wind density and $\Sigma_{\rm j}=\pi r^{2}_{\rm h}\theta^{2}_{\rm j}$ is the jet cross-section at the location of the head with $\theta_{\rm j}$ being the jet head opening angle. $L_{\rm j,iso}=2L_{\rm j}/\theta_{\rm j}^{2}$ is the isotropic equivalent luminosity and $\dot{M}_{\rm w}=4\pi r^{2}_{\rm h}\rho_{\rm w}\beta_{\rm w}c$ is disk wind mass-loss rate.
Figure 1: Evolution of $\theta_{\rm j}/\theta_{0}$ as a function of the jet efficiency $\eta$ for various initial jet opening angles $\theta_{0}$.
@@
$\begin{split}\theta_{\rm j}=\left\{\begin{array}[]{ll}\tilde{L}^{1/4}\theta_{0}^{2}=(\frac{2\eta\beta_{\rm w}}{f_{\rm w}})^{1/6}\theta_{0}^{4/3},&\tilde{L}<1\\ \tilde{L}^{3/4}\theta_{0}^{2}=(\frac{2\eta\beta_{\rm w}}{f_{\rm w}})^{3/10}\theta_{0}^{4/5},&1<\tilde{L}<\theta_{0}^{-4/3}\\ \theta_{0}&\tilde{L}>\theta_{0}^{-4/3}\end{array}\right.\end{split}$ (7)
where $\eta=L_{\rm j}/\dot{M}c^{2}$ is the jet efficiency. Fig.1 shows the evolution of $\theta_{\rm j}$ as a function of the jet efficiency $\eta$. It indicates that for low jet efficiency, the jet head opening angle is significantly lower than the initial opening angle of the jet as the result of the collimation. The parameter $\tilde{L}$ also indicates the velocity of jet head. For $\tilde{L}<1$, the jet head is non-relativistic and vice versa.
@@
$L_{j}\geq\frac{1}{2}\dot{M}f_{\rm w}c^{2}\beta_{\rm w}\pi^{2}{\rm sin}^{2}\theta_{\rm LS}.$ (8)
@@
Figure 2: Break-out condition in terms of the jet efficiency $L_{\rm j}/\dot{M}c^{2}$ and misaligned angle $\theta_{\rm LS}$ for the parameter values of $\rm\theta_{0}=0.33$, $f_{\rm w}=0.5$, $\beta_{\rm w}=0.1$. The gray solid line shows the jet break-out criterion. The darker regions indicate break-out jet, whereas the lighter regions indicate choked jet. The yellow regions, the red regions and the gray region represent collimated jets with non-relativistic jet head, collimated jets with relativistic jet head and uncollimated jets, respectively. The horizontal black solid line shows the case of Swift J1644+57 with a jet efficiency of $\eta=0.023$.
@@
Figure 3: Break-out condition in terms of the dimensionless term $L_{\rm j,iso}/\beta_{\rm w}\dot{M}_{\rm w}c^{2}$ and misaligned angle $\theta_{\rm LS}$ using the same parameters as those in Fig. 2. The solid gray line represents the break-out criterion when the the collimation caused by the cocoon pressure is taken in into account, while the blue dot-dashed line shows the break-out condition without taking into account the collimation effect (i..e, assuming that jets are conical). The meanings of the regions with different colors are the same as that in Fig.2.
@@
$\begin{split}\frac{L_{\rm j,iso}}{\beta_{\rm w}\dot{M}_{\rm w}c^{2}}>\left\{\begin{array}[]{ll}(\beta_{\rm w}^{1/2}\pi^{2}{\rm sin}^{2}\theta_{\rm LS}\theta_{0}^{-4})^{2/3}/\beta_{\rm w},&\tilde{L}<1\\ (\beta_{\rm w}^{-1/2}\pi^{2}{\rm sin}^{2}\theta_{\rm LS}\theta_{0}^{-4})^{2/5}/\beta_{\rm w},&1<\tilde{L}<\theta_{0}^{-4/3}\\ \pi^{2}\rm sin^{2}\theta_{\rm LS}\theta_{0}^{-2},&\tilde{L}>\theta_{0}^{-4/3}\end{array}\right.\end{split}$ (9)
@@ 3 High-energy neutrinos from choked precessing jets
The shocks in choked jets will accelerate cosmic rays to ultra-high energies (Farrar & Gruzinov, 2009; Wang et al., 2011; Wang & Liu, 2016) and further produce high-energy neutrinos (Wang & Liu, 2016; Mukhopadhyay et al., 2023; Lu et al., 2023). Internal shocks may occur due to the internal collisions within the jets, resulted from the inhomogeneity in the jet velocity. Temporal variability with $\delta t\approx 100s$ has been seen in the X-ray emission of the jetted TDE Swift J1644+57, which is thought to be generated by internal shocks (Burrows et al., 2011). We take a fiducial physical value $\Gamma\sim 1/\theta_{0}$ for the jet Lorentz factor, and hence the collision occurs at $R_{\rm int}\approx\Gamma^{2}c\delta t=3\times 10^{14}{\rm cm}\Gamma_{1}^{2}$. Internal shocks that propagate into the low-density jets are collisionless, although they locate inside the optically thick envelope (Wang & Liu, 2016).
One important process for neutrino production in choked jets is the interaction between the cosmic rays and the thermal photons generated in the wind envelope (i.e., the $p\gamma$ process). To calculate the neutrino flux, we need to know the the density of thermal photons in the jet. Since the jet size is much smaller than that of the wind envelope, the density of thermal photons around the jet should be roughly equal to the density of thermal photons in the envelope at the same radius. We use the same treatment of thermal photons as that in Zheng et al. (2023). The photosphere temperature can be obtained from observations of TDEs. Then from the energy transferring equation, we can derive the temperature inside the envelope, which follows $T\propto r^{-(n+1)/4}$ for $r\ll R_{\rm ph}$ (Roth et al., 2016), where $R_{\rm ph}$ is the photosphere radius. For the wind density, we have $n=2$ and $T\propto r^{-3/4}$. The typical energy of protons interacting with thermal photons with temperature $T=10^{5}-10^{6}{\rm K}$ is $\epsilon_{\rm p}=0.1{\rm GeV}^{2}/3kT=0.6-6{\rm PeV}$ and the neutrino energy is correspondingly $\epsilon_{\nu}=0.05\epsilon_{\rm p}=30-300{\rm TeV}$, consistent with the measured energy of three neutrinos possibly associated with TDEs.
Considering that the duty cycle of the neutrino radiation cone for a precessing jet is $\chi_{\rm duty}=(1/\Gamma)/\pi{\rm sin}\theta_{\rm LS}$, the fluence of neutrinos from this precessing jet is
@@
where $t_{d}$ is the neutrino trigger time, $L_{\rm p}=\varepsilon_{\rm p}L_{\rm j}\Gamma^{2}$ is the isotropic proton luminosity (assuming a neutrino radiation cone of $1/\Gamma$), $\varepsilon_{\rm p}$ is the fraction of the jet energy converted into relativistic protons and $\Gamma$ is the bulk Lorentz factor of the jet. $f_{\rm p\gamma}$ is the fraction of energy loss by the $p\gamma$ process, which can be obtained following Zheng et al. (2023). Assuming that the jet power scales with the accretion rate, the jet luminosity evolves with time as $L_{\rm j}\propto(t/\tau)^{-\frac{5}{3}}$ after the peak (Krolik & Piran, 2012; Piran et al., 2015), where $\tau$ is the characteristic fallback timescale, which is roughly the orbital period of the most bound debris (Gezari, 2021), i.e.,
@@
with $M_{\rm BH}$ being the mass of SMBH and $M_{\star}$ being the mass of the disrupted star. We take the minimum energy of protons as $\epsilon_{\rm p,min}=\Gamma m_{\rm p}c^{2}$ and assume equipartition factors $\varepsilon_{\rm B}=0.1$ and $\varepsilon_{\rm p}=0.2$ in the calculation. Then we can obtain the cumulative muon neutrino number between 10 TeV and 1PeV by using
@@
Figure 5: Color map of the expected neutrino numbers for three TDEs (AT2019dsg, AT2019fdr, AT2019aalc) as a function of the jet luminosity $L_{\rm j}$ and initial opening angle $\theta_{0}$ with a fixed misaligned angle $\theta_{\rm LS}=\pi/4$. The white dashed line represent the minimal required neutrino numbers inferred from observations. The black vertical line represents the maximum jet luminosity limited by the break-out criterion represented as $L_{\rm BO}$, while the white solid line represents the maximum jet luminosity limited by the wind power ($\frac{1}{2}\dot{M}_{\rm w}v_{\rm w}^{2}$).
Figure 6: Color map of the expected neutrino numbers for three TDEs (AT2019dsg, AT2019fdr, AT2019aalc) as a function of the jet luminosity $L_{\rm j}$ and misaligned angle $\theta_{\rm LS}$ with a fixed initial jet half-opening angle of $\theta_{0}=0.1$. The meanings of each lines are the same as that in Fig.5.
We calculate the expected neutrino number from precesing jets as the function of the misaligned angle and jet power, as shown in Fig.4. We take fiducial values of $\Gamma=3$, $\theta_{0}=0.33$ and $\tau=41{\rm day}$ (corresponding to a SMBH with $10^{6}M_{\odot}$) for the jets. We take a reference value ($D={\rm 220Mpc}$) for the TDE distance, equal to the distance of AT2019dsg. As expected, the neutrino flux increases as the jet power increases. The neutrino flux also increases as the misaligned angle decreases, because the duty cycle that the jet sweeps across increases as the misaligned angle decreases. The black solid line represents the break-out condition given by Eq.(7). The maximum neutrino number that precessing choked TDEs can produce is $N_{\nu}\sim 0.1$. The black dot-dashed line shows the case of TDEs like Swift J1644+57, which has a high luminosity of $10^{46}\rm erg~{}s^{-1}$. For this case, the maximum neutrino number produced by the choked jet is about $N_{\nu}\sim 0.1$.

### v3

@@ 2 Jet propagation and the break-out condition for precessing jets
After the stellar disruption, an accretion disk is formed in the orbital plane of the star whose normal is generally misaligned by an angle $\theta_{\rm LS}$ with respect to the SMBH spin axis. The accretion disk undergoes Lense-Thirring precession around the black hole spin axis and the jet is aligned with the instantaneous disk angular momentum (Teboul & Metzger, 2023; Lu et al., 2023). Matter falls back to the disk at a highly super-Eddington accretion rate $\dot{M}$, driving a disk wind of velocity $\beta_{\rm w}=v_{\rm w}/c$ and mass-loss rate $\dot{M}_{\rm w}=f_{\rm w}\dot{M}$, where $f_{\rm w}<1$ is the fraction of the mass fallback rate lost in the wind. Metzger & Stone (2016) argued that only a small fraction the returning debris ends up reaching the inner circular disk and accreting onto the SMBH, with the vast majority becoming unbound in a wind outflow. So we consider $f_{\rm w}=0.5$ in our calculation. We choose $\beta_{\rm w}=0.1$ as a typical wind velocity. Due to the rapid precession, these disk winds will inevitably interact and collide with each other on large scales, leading to a quasi-spherical outflow.
We consider a relativistic jet with a kinetic luminosity $L_{\rm j}$, an initial half-opening angle $\theta_{0}$ and a bulk Lorentz factor $\Gamma=10\gtrsim 1/\theta_{0}$. As the jet advances in the quasi-spherical wind envelope, the jet drives a bow shock ahead of it. The jet is capped by a termination shock, and a reverse shock propagates back into the jet, where the jet is decelerated and heated. The hot head material spills sideways, forming a cocoon that engulfs the jet and collimate it. The velocity of the jet head is determined by the balance of pressure between ambient gas and the jet head (Matzner, 2003). In a moving wind envelope, we can express the velocity in the frame of the wind envelope by a coordinate transformation (e.g., Ioka & Nakamura, 2018; Hamidani et al., 2020; Gottlieb & Nakar, 2022; Lu et al., 2023):
@@
$\tilde{L}=\frac{L_{\rm j}}{\Sigma_{\rm j}\rho_{\rm w}(r_{\rm h})c^{3}}=\frac{2L_{\rm j,iso}\beta_{\rm w}}{\dot{M}_{\rm w}c^{2}}.$ (2)
Here $\beta_{\rm h}$ is the jet head velocity in unit of the speed of light (measured in the lab frame), $r_{\rm h}$ is the radius of the jet head, $\rho_{\rm w}$ is wind density and $\Sigma_{\rm j}=\pi r^{2}_{\rm h}\theta^{2}_{\rm j}$ is the jet cross-section at the location of the head with $\theta_{\rm j}$ being the jet head opening angle [^1] The jet head opening angle is defined as $\theta_{\rm j}=r_{\rm j}/r_{\rm h}$ (regardless of the jet collimation), where $r_{\rm j}$ is the radius of jet head cross-section and $r_{\rm h}=\beta_{\rm h}ct$ is the height of the jet head.. $L_{\rm j,iso}=2L_{\rm j}/\theta_{\rm j}^{2}$ is the isotropic equivalent luminosity and $\dot{M}_{\rm w}=4\pi r^{2}_{\rm h}\rho_{\rm w}\beta_{\rm w}c$ is disk wind mass-loss rate.
Figure 1: The evolution of $\theta_{\rm j}$ adopting Eq.(7). Top: $\theta_{\rm j}/\theta_{0}$ as a function of $\tilde{L}$ for various initial jet opening angles $\theta_{0}$. Bottom: $\theta_{\rm j}/\theta_{0}$ as a function of the jet efficiency $\eta$ for various initial jet opening angles $\theta_{0}$.
@@
$\begin{split}\theta_{\rm j}=\left\{\begin{array}[]{ll}\tilde{L}^{1/4}\theta_{0}^{2}=(\frac{4\eta\beta_{\rm w}}{f_{\rm w}})^{1/6}\theta_{0}^{4/3},&\tilde{L}<1\\ \tilde{L}^{3/4}\theta_{0}^{2}=(\frac{4\eta\beta_{\rm w}}{f_{\rm w}})^{3/10}\theta_{0}^{4/5},&1<\tilde{L}<\theta_{0}^{-4/3}\\ \theta_{0}&\tilde{L}>\theta_{0}^{-4/3}\end{array}\right.\end{split}$ (7)
where $\eta=L_{\rm j}/\dot{M}c^{2}$ is the jet efficiency. Fig.1 shows the evolution of $\theta_{\rm j}$ as a function of the dimensionless jet luminosity $\tilde{L}$ and the jet efficiency $\eta$. The upper panel shows that the dimensionless parameter $\tilde{L}$ can determine the jet head opening angle and velocity of jet head. For $\tilde{L}<1$, the jet head is non-relativistic and vice versa. The lower panel indicates that for a low jet efficiency, the jet head opening angle is significantly lower than the initial opening angle of the jet as the result of the collimation.
@@
$L_{\rm j}\geq\frac{1}{2}\dot{M}f_{\rm w}c^{2}\beta_{\rm w}\pi^{2}{\rm sin}^{2}\theta_{\rm LS}.$ (8)
@@
Figure 2: Break-out condition in terms of the jet efficiency $L_{\rm j}/\dot{M}c^{2}$ and misaligned angle $\theta_{\rm LS}$ for the parameter values of $\rm\theta_{0}=0.33$, $f_{\rm w}=0.5$, $\beta_{\rm w}=0.1$. The gray solid line shows the jet break-out criterion. The darker regions indicate break-out jet, whereas the lighter regions indicate choked jet. The yellow regions, the red regions and the gray region represent collimated jets with non-relativistic jet head, collimated jets with relativistic jet head and uncollimated jets, respectively. The horizontal black solid line shows the case of Swift J1644+57 with a jet efficiency of $\eta=0.023$. Note that we simply just use the region larger than unity to denote the figure labels.
@@
Figure 3: Break-out condition in terms of the dimensionless term $L_{\rm j,iso}/\beta_{\rm w}\dot{M}_{\rm w}c^{2}$ and misaligned angle $\theta_{\rm LS}$ using the same parameters as those in Fig. 2. The solid gray line represents the break-out criterion when the the collimation caused by the cocoon pressure is taken in into account, while the blue dot-dashed line shows the break-out condition without taking into account the collimation effect (i.e., assuming that jets are conical). The meanings of the regions with different colors are the same as that in Fig.2.
@@
$\begin{split}\frac{L_{\rm j,iso}}{\beta_{\rm w}\dot{M}_{\rm w}c^{2}}>\left\{\begin{array}[]{ll}(2^{-1/2}\beta_{\rm w}^{1/2}\pi^{2}{\rm sin}^{2}\theta_{\rm LS}\theta_{0}^{-4})^{2/3}/\beta_{\rm w},&\tilde{L}<1\\ (2^{-3/2}\beta_{\rm w}^{-1/2}\pi^{2}{\rm sin}^{2}\theta_{\rm LS}\theta_{0}^{-4})^{2/5}/\beta_{\rm w},&1<\tilde{L}<\theta_{0}^{-4/3}\\ \pi^{2}\rm sin^{2}\theta_{\rm LS}\theta_{0}^{-2},&\tilde{L}>\theta_{0}^{-4/3}\end{array}\right.\end{split}$ (9)
@@ 3 High-energy neutrinos from choked precessing jets
The shocks in choked jets will accelerate cosmic rays to ultra-high energies (Farrar & Gruzinov, 2009; Wang et al., 2011; Wang & Liu, 2016) and further produce high-energy neutrinos (Wang & Liu, 2016; Mukhopadhyay et al., 2023; Lu et al., 2023). Internal shocks may occur due to the internal collisions within the jets, resulted from the inhomogeneity in the jet velocity. Temporal variability with $\delta t\approx 100s$ has been seen in the X-ray emission of the jetted TDE Swift J1644+57, which is thought to be generated by internal shocks (Burrows et al., 2011). We take a fiducial physical value $\Gamma\gtrsim 1/\theta_{0}$ for the jet Lorentz factor, and hence the collision occurs at $R_{\rm int}\approx\Gamma^{2}c\delta t=3\times 10^{14}{\rm cm}\Gamma_{1}^{2}$. Internal shocks that propagate into the low-density jets are collisionless, although they locate inside the optically thick envelope (Wang & Liu, 2016).
One important process for neutrino production in choked jets is the interaction between the cosmic rays and the thermal photons generated in the wind envelope (i.e., the $p\gamma$ process). The cooling timescale of $p\gamma$ process is
@@
where $\gamma_{p}$ is Lorentz factors of protons, ${dn}/{d\epsilon}$ is the number density of seed photons, $\epsilon_{\text{th}}\simeq 145{\rm MeV}$ is the threshold energy of $p\gamma$ interaction, $\sigma_{p\gamma}$ is cross-section of $p\gamma$ interaction and $\kappa_{p\gamma}$ is inelasticity. To calculate the neutrino flux, we need to know the the density of thermal photons in the jet. Since the jet size is much smaller than that of the wind envelope, the density of thermal photons around the jet should be roughly equal to the density of thermal photons in the envelope at the same radius. We use the same treatment of thermal photons as that in Zheng et al. (2023). The photospheric temperature can be obtained from observations of TDEs. When the diffusion timescale is shorter than the dynamic timescale, photons escape from the optically thick envelope through diffusion. Then from the energy transferring equation, we can derive the temperature inside the envelope, which follows $T\propto r^{-(n+1)/4}$ for $r\ll R_{\rm ph}$ (Roth et al., 2016), where $R_{\rm ph}$ is the photosphere radius. For the quasi-isotropic wind , we have $n=2$ and $T\propto r^{-3/4}$. On the other hand, when the diffusion timescale exceeds the dynamic timescale, photons become trapped within the fluid, leading to that the radiation temperature is determined only by the adiabatic cooling. In this case, the temperature of wind evolves as $T\propto r^{-2/3}$ for quasi-isotropic wind with $n=2$ (Piro & Lu, 2020; Matsumoto & Piran, 2021a). The typical energy of protons interacting with thermal photons with temperature $T=10^{5}-10^{6}{\rm K}$ is $\epsilon_{\rm p}=0.1{\rm GeV}^{2}/3kT=0.6-6{\rm PeV}$ and the neutrino energy is correspondingly $\epsilon_{\nu}=0.05\epsilon_{\rm p}=30-300{\rm TeV}$, consistent with the measured energy of three neutrinos possibly associated with TDEs.
Considering that the duty cycle of the neutrino radiation cone for a precessing jet is $\chi_{\rm duty}=\theta_{b}/\pi{\rm sin}\theta_{\rm LS}$, where $\theta_{b}={\rm max}\{\theta_{\rm j},1/\Gamma\}$ is the half opening angle of neutrino radiation cone. The fluence of neutrinos from this precessing jet is
[…]
where $t_{d}$ is the neutrino trigger time, $L_{\rm p}=\varepsilon_{\rm p}L_{\rm iso}$ is the isotropic proton luminosity which is defined as $L_{\rm iso}=2L_{\rm j}/\theta_{b}^{2}$, $\varepsilon_{\rm p}$ is the fraction of the jet energy converted into relativistic protons and $\Gamma$ is the bulk Lorentz factor of the jet. $f_{\rm p\gamma}$ is the fraction of energy loss by the $p\gamma$ process, which can be expressed as $t_{\rm p\gamma}^{-1}/(t_{\rm p\gamma}^{-1}+t_{\rm dyn}^{-1})$, where $t_{\rm dyn}\approx R_{\rm int}/c$ is dynamical timescale of the jet. In our calculation $f_{\rm p\gamma}\simeq 1$ is obtained. Assuming that the jet power scales with the accretion rate, the jet luminosity evolves with time as $L_{\rm j}\propto(t/\tau)^{-\frac{5}{3}}$ after the peak (Krolik & Piran, 2012; Piran et al., 2015), where $\tau$ is the characteristic fallback timescale, which is roughly the orbital period of the most bound debris (Gezari, 2021), i.e.,
@@
with $M_{\rm BH}$ being the mass of SMBH and $M_{\star}$ being the mass of the disrupted star. The maximum energy of protons is obtained by equating the acceleration time $t_{\rm acc}=\gamma_{\rm p}m_{\rm p}c/\eta_{\rm acc}eB$ to the cooling time $t_{\rm p\gamma}$, where $\eta_{\rm acc}=0.1$ is acceleration efficiency, and magnetic field is given by $B=\sqrt{4\varepsilon_{\rm B}L_{\rm j}\Gamma^{2}/R_{\rm int}^{2}c}$. We take the minimum energy of protons as $\epsilon_{\rm p,min}=\Gamma m_{\rm p}c^{2}$ and assume equipartition factors $\varepsilon_{\rm B}=0.1$ and $\varepsilon_{\rm p}=0.2$ in the calculation. Then we can obtain the cumulative muon neutrino number between 10 TeV and 1PeV by using
@@
Figure 5: Color map of the expected neutrino numbers for three TDEs (AT2019dsg, AT2019fdr, AT2019aalc) as a function of the jet efficiency $\eta$ and initial opening angle $\theta_{0}$ with a fixed misaligned angle $\theta_{\rm LS}=\pi/4$. The white dashed line represent the minimal required neutrino numbers inferred from observations. The black vertical line represents the maximum jet efficiency limited by the break-out criterion, while the white solid line represents the critical jet efficiency $\eta_{c}=0.0025$.
Figure 6: Color map of the expected neutrino numbers for three TDEs (AT2019dsg, AT2019fdr, AT2019aalc) as a function of the jet efficiency $\eta$ and misaligned angle $\theta_{\rm LS}$ with initial jet half-opening angle of $\theta_{0}=0.33$ and $\Gamma=10$. The meanings of each lines are the same as that in Fig.5.
We calculate the expected neutrino number from precesing jets as the function of the misaligned angle and jet power, as shown in Fig.4. We take fiducial values of $\Gamma=10$, $\theta_{0}=0.33$ and $\tau=41{\rm day}$ (corresponding to a SMBH with $10^{6}M_{\odot}$) for the jets. We take a reference value ($D={\rm 220Mpc}$) for the TDE distance, equal to the distance of AT2019dsg. As expected, the neutrino flux increases as the jet power increases. The neutrino flux also increases as the misaligned angle decreases, because the duty cycle that the jet sweeps across increases as the misaligned angle decreases. The black solid line represents the break-out condition given by Eq.(8). The maximum neutrino number that precessing choked TDEs can produce is $N_{\nu}\sim 0.1$. The black dot-dashed line shows the case of TDEs like Swift J1644+57, which has a high luminosity of $10^{46}\rm erg~{}s^{-1}$. For this case, the maximum neutrino number produced by the choked jet is about $N_{\nu}\sim 0.1$.